Problem Page A textbook store sold a combined total of 203 history and sociology textbooks in a week. The number of sociology textbooks sold was 47 less than the number of history textbooks sold. How many textbooks of each type were sold?
Answer by addingup(3677)

h + s = 203 subtract s from both sides:
h = 203 - s
s = h - 47 so substitute for s in the first equation above:
h + h - 47 203
2h = 250
h = 125 they sold 125 history books
and sociology?:
s = h - 47 = 125 - 47 = 78 books